Six vitamin and four sugar tablets identical in appearance are in a box. One tablet is taken at random and given to Person A. Then a second tablet is selected and given to Person B. What is the probability that Type answers in NON-SIMPLIFIED FRACTIONS! Person A was given a sugar tablet? Person B was given a vitamin tablet given that Person A was given a sugar tablet? a. c. neither was given a vitamin tablet? d. both were given a vitamin tablet? e. exactly one person was given a vitamin tablet? Person A was given a sugar tablet and Person B was given a sugar tablet? f, Person A was given a vitamin tablet and Person B was given a sugar tablet? g.
Answer #1

a)P(person A given a sugar tablet)=4/10=2/5

b)P(person B vitamin|A Sugar)=6/9=2/3

c)P(neither Vitamin)=(4/10)*(3/9)=(2/5)*(1/3)=2/15

d)P(both Vitamin)=(6/10)*(5/9)=30/90=1/3

e)exactly one Vitamin =P(A VItamin and B Sugar)+P(B VItamin and A Sugar)

=(6/10)*(4/9)+(4/10)*(6/9)=48/90=24/45=8/15

f)P(both Sugar)=(4/10)*(3/9)=12/90=2/15

g)P(A VItamin and B Sugar)=(6/10)*(4/9)=24/90=4/15